OpenStudy (anonymous):

I'm using u substitution to do this integral. I understand how to do the integral, but I don't know how adjust the limits of the integral. So.... Int from pi/2 to 3pi/4 of sin(^5)x*cos(^3)xdx =Int from pi/2 to 3pi/4 of sin(^5)x*(1-sin(^2)x)*cosxdx u=sinx du=cosxdx ....so now how do i adjust the limits (from pi/2 to 3pi/4) of the integral

OpenStudy (anonymous):

plug in the original limits in whatever you used for u. if you used sinx, plug in the limits in x and you have your new limits

OpenStudy (mathmate):

Think of the original limits as from x=pi/2 to 3pi/4. The new limits are now: u=sin(pi/2) to u=sin(3pi/4). since the variable of integration is now u.

OpenStudy (anonymous):

I see...thank you

OpenStudy (mathmate):

You're welcome!